OpenStudy (anonymous):

Nakim simplified 3 times the square root of 2x plus x times the square root of 8x minus 5 times the square root of 18x and got -10x times the square root of 2x for an answer. Part 1: Using complete sentences, explain what Nakim did wrong. Part 2: Show all your work to simplify the expression. I dont understand this one at all! please help

OpenStudy (cwtan):

This question has been asked by others http://openstudy.com/updates/4fc653d5e4b022f1e12dcb0f

OpenStudy (cwtan):

\[\sqrt {a^2b}=a\sqrt {b}\] \[5\sqrt {18x}=5\sqrt {9\times 2x}=5(3\sqrt {2x})=15\sqrt {2x}\] \[x\sqrt {8x}=x\sqrt {4\times 2x}=2x\sqrt {2x}\] so \[3\sqrt {2x}+x\sqrt{2x}-5\sqrt {18x}\] \[=3\sqrt {2x}+2x\sqrt {2x}-15\sqrt {2x}\] \[=(3-15+2x)\sqrt {2x}\] \[=(-12+2x)\sqrt{2x}\]

OpenStudy (cwtan):

there was a small typing error.....